OpenStudy (anonymous):

factor the trinomial: w^3-2w^2-8w=

OpenStudy (dumbcow):

=w(w-4)(w+2)

OpenStudy (anonymous):

= w ( w^2 - 2w - 8) = w ( w+2) (w-4)